The weight of a body of area 20 Sq cm and height 10 cm in air and water is 9.8 N and 7.84 N respectively. Calculate the density of the material of the body.

Grade:10

1 Answers

vijay kale
6 years ago
Mass=9.8N/9.8ms^-2=1kg=1000gVolume=base area×h=20×10=200 cubic cmDensity=mass/volume=1000g/200 cubic cm=5gm per cubic cm